No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>Search

Reminder

To have a better experience, please upgrade your IE browser.

upgrade
Knowledge Base

MA5633 R17 - Issue with DVB-C output power level

Publication Date:  2018-09-20  |   Views:  574  |   Downloads:  0  |   Author:  f84102783  |   Document ID:  EKB1001869026

Contents

Issue Description

I have a MA5633 R17 in GPoN mode here.
We had to replace this morning because output 2/3 of the levels can not be adjusted (digiTV).
The level stays about 10 dB below the nominal level and does not make any more faltering.
Output 0/1 are still ok and Docsis are ok.

Solution

We have two groups, group 1 for output 0/1, group 2 for output 2/3.

 

(1) The analysis of signal measured.

<1> configuration of EQ and ATT


 

<2> Signal of group 1


<3> Signal of group 2


 

We can find the level of DOCSIS channel of group 2 is 5db, below the level of group 1 and the attenuation of group 1 is 3db the attenuation of group 2 is 0db.

So, the DOCSIS channel output of group 2 is 8db bellow the output of group1.

So, this DOCSIS channel is not ok too.

The output of group is abnormal for all frequency, but for high frequency the output is more bad.

 

(2) The analysis of rf-power-monitor.

In order to confirm our conclusion, we shutdown optical- receiver and reconfigure the DOCSIS channel from frequency 250MHZ to 826MHZ.

<1> The new configuration of DOCSIS channel


<2> The new configuration of EQ and ATT

We configure the same ATT and EQ of group 1 and group 2, and the operation ATT and EQ is 3db and 18db too.


 

<3> The output level of group 1 and group 2

We measured the output level of group 1 and group 2 with command bellow, and create the output graph.


Output graph of 2 group:


The EQ of group is abnormal, but the configuration and status of EQ is OK, so, we confirm hardware broken.

 

Conclusion:

We confirm the hardware of this device is damaged. And the trouble spots is EQ chip or the circuit around EQ chip of group 2.